# MultiChartQA: Benchmarking Vision-Language Models on Multi-Chart Problems

### v1

@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have received significant attention due to their impressive capabilities across a broad set of vision-language tasks, such as image captioning Agrawal et al. (2019); Wang et al. (2023), multimodal dialogue systems Saha et al. (2018); Yin et al. (2024), and visual question answering (VQA) Schwenk et al. (2022); Li et al. (2024b); Jia et al. (2024a). Among their varied applications, the comprehension of visual chart data, such as bar and line charts, holds particular significance Han et al. (2023); Zhang et al. (2024a). Such data is prevalent in real-world scenarios, including academic papers and analytical reports, which makes understanding and reasoning over these charts an essential skill for MLLMs. As a result, several efforts have been made in evaluating and improving model performance on chart-related tasks Masry et al. (2022); Xu et al. (2023); Masry et al. (2024b).
@@
To bridge this gap, we introduce MultiChartQA, a newly collected benchmark designed to evaluate the capabilities of MLLMs on multi-chart scenarios. Our benchmark draws from a diverse set of public resources containing multi-chart articles, including Arxiv, OECD, OWID, and the Pew Research Center. Using the charts collected from these sources, we design four different question types, each paired with multiple charts from the same article as input:
@@
We evaluate the performance of 16 mainstream MLLMs on MultiChartQA. The results reveal a significant limitation in MLLMs’ ability to process multiple charts, with performance trailing behind humans by a large margin. Furthermore, there is a notable gap between the performance of proprietary models and their open-source counterparts. Further analysis conducted on our benchmark highlights the benefits of using chain-of-thought reasoning, as well as the models’ reliance on chart references for information retrieval. An in-depth error analysis reveals that MLLMs still struggle with visual perception and multi-step reasoning, especially when color or spatial constraints are involved. We believe that MultiChartQA will illuminate the evaluation of multi-chart capabilities in MLLMs and foster advancements in this field in the near future.
@@ 3 MultiChartQA
MultiChartQA is an extensive and demanding benchmark that features real-world charts and complex reasoning questions. To cover charts with various topics and a diverse range of styles, we curate chart sets from a variety of real-world sources (§3.1). Based on four core competencies in chart comprehension, we design four categories of question-answer pairs (§3.2). We detail the evaluation metrics in §3.3. Ultimately, the MultiChartQA benchmark comprises 655 charts and 944 questions.
@@
- Direct Question 212 (22.4%)
- Parallel Questions 315 (33.4%)
- Comparative Reasoning 322 (34.1%)
- Sequential Reasoning 95 (10.1%)
@@
-Multiple-choice 60 (6.4%)
- Open-ended 884 (93.6%)
Unique charts 655
Multi-chart sets 236
Average charts 2.8
Source
ArXiv/OECD/OWID/Pew 64/58/58/56
@@
Unique questions 761
Unique tokens 2,593
Maximum/Average length 162/39.9
@@
Unique answers 518
Unique tokens 858
Maximum/Average length 55/3.4
@@ 3.1 Chart Collection
We aim to employ chart sets with semantic connection and topic consistency for MultiChartQA. The charts within the same article or paper naturally fulfill this criterion. We thus identify four websites that feature such multi-chart content: ArXiv, OECD, OWID, and Pew, and collect chart sets from these sources.
ArXiv [^1] https://arxiv.org/ is an open-access repository that hosts preprints of research papers across various scientific disciplines. We extracted images from 1,500 publications using the arXiv API [^2] https://info.arxiv.org/help/api/index.html. Since research papers often contain a variety of images beyond charts, such as natural images or diagrams, we employ an off-the-shelf MLLM, LLaVA-1.5-13B Liu et al. (2024a), in a zero-shot prompting approach to classify and filter out non-chart images. We consider multiple charts from the same paper as a chart set. To ensure the charts in one set have strong semantic correlations, we manually review the sets and retain those that share one or more elements, e.g., same entities or categories. In total, we collect 64 sets of charts from Arxiv. OECD [^3] https://www.oecd.org/ is a platform that hosts reports focused on data and analysis. Each report typically includes multiple charts illustrating data across different time spots. We treat the charts from the same report as a semantically related chart set. In total, we collect 58 sets of multi-chart samples from the OECD. OWID [^4] https://ourworldindata.org/ is an online scientific publication platform that features articles with extensive chart demonstrations across various fields. Similar to the OECD, we treat charts within a single report as a chart set. In total, we collect 58 sets of charts from OWID. Pew [^5] https://www.pewresearch.org/ publishes reports including multiple charts as rich visual presentations. Similar to OECD, we consider charts within each report as a chart set, resulting in the collection of 56 multi-chart sets. Overall, 236 multi-chart sets are collected, each containing 2-3 semantically related charts. Examples of charts from different sources are demonstrated in Appendix G.
@@ 3.2 Question Design
The direct questions evaluate whether the model can accurately identify the relevant chart to answer questions accordingly. We present multiple charts and use terms like “In the second chart” to explicitly specify which chart the answer should pertain to.
The design of the questions focuses on two types of chart information: structural information and content information. The former involves details about the chart’s structure, such as the values on the axes, the layout of subplots, and the labels in the legend Wang et al. (2024c). These details are essential for comprehending the overall architecture of the chart and interpreting its organizational form. The latter focuses on the chart’s content, such as extracting specific numbers or phrases based on the given clues. This type of task requires the model to understand the semantics of the chart content to perform precise reasoning and information analysis.
@@
Claude-3.5-Sonnet 75.30 85.71 77.22 86.75 84.56 75.78 71.67 70.42 52.63
GPT-4o 69.58 78.20 78.48 77.99 80.28 69.53 75.00 58.10 37.89
Gemini-1.5-Pro 59.68 72.93 63.29 72.86 70.18 54.30 59.44 49.65 37.89
GPT-4o-mini 51.18 60.90 65.82 51.28 53.36 48.05 52.78 46.83 30.53
@@
MiniCPM-V2.6 45.62 66.17 62.03 52.99 59.48 39.84 26.67 46.13 24.21
InternVL2-26B 44.88 66.17 55.70 47.86 57.65 42.19 32.78 40.49 22.11
InternVL-V1.5 39.16 57.89 48.10 40.81 51.68 33.98 28.33 37.32 20.00
Qwen2-VL-7B 39.02 60.15 58.23 45.51 48.01 41.80 16.11 41.55 13.68
LLaVA-OV-7B 32.22 45.86 51.90 40.38 41.90 34.77 12.22 34.15 10.53
MiniCPM-V2.5 31.51 48.87 20.25 38.03 44.34 28.12 21.11 30.63 22.11
DeepSeek-VL-7B 24.91 41.35 27.85 25.64 28.13 24.22 12.22 32.75 8.42
LLaVA-V1.6-7B 22.86 27.07 25.32 25.21 35.93 26.56 8.89 23.94 17.89
Idefics3-8B 10.24 10.53 12.66 10.47 17.43 15.62 2.22 13.03 3.16
Idefics2-8B 8.83 6.77 16.46 11.75 13.91 8.59 5.56 8.45 4.21
@@
MatCha 3.60 3.76 6.33 2.78 3.06 1.95 1.67 6.34 4.21
ChartInstruct 3.04 6.77 1.27 5.34 4.13 2.34 0.00 3.52 2.11
Table 2: Evaluation results on MultiChartQA. Bold values indicate the best performance within each category: closed-source or open-source models. Question categories and model names are abbreviated due to space limits. Struct.: Structure, Cont.: Content, Seq.: Sequential, LLaVA-OV-7B: LLaVA-OneVision-7B, ChartInstruct: ChartInstruct-LLama2
@@ 3.3 Evaluation Metric
We add explicit instructions for each question to control the output format, enabling efficient extraction of answers. The model outputs are parsed and answers are extracted using scripts [^6] https://github.com/MMMU-Benchmark/MMMU.. Then we use string-matching to check if the extracted answer matches the ground truth. We adopt accuracy as the evaluation metric. For questions that consist of multiple sub-questions, i.e., parallel questions, we calculate accuracy based on the proportion of correctly answered sub-questions, ensuring a fair and fine-grained evaluation.
@@ 4.1 Experimental Setup
We evaluate 16 mainstream MLLMs on our benchmark, comprising 4 closed-source models, 10 open-source general-purpose models, and 2 chart-domain specialized models. The closed-source models include Claude-3.5-Sonnet Anthropic (2024), GPT-4o OpenAI (2024), GPT-4o-mini OpenAI (2024), and Gemini-1.5-Pro Reid et al. (2024). Among the open-source general-purpose models, we evaluate MiniCPM-V2.6 Yao et al. (2024), InternVL2-26B Chen et al. (2024), InternVL-V1.5 Chen et al. (2024), Qwen2-VL-7B Wang et al. (2024b), LLaVA-OneVision-7B Li et al. (2024a), LLaVA-V1.6-7B Liu et al. (2024b), MiniCPM-V2.5 Yao et al. (2024), DeepSeek-VL-7B Lu et al. (2024), Idefics3-8B Laurençon et al. (2024a), and Idefics2-8B Laurençon et al. (2024b). For the chart- model, we evaluate MatCha Liu et al. (2022) and ChartInstruct-LLama2 Masry et al. (2024a). For reference, we include human performance and random choice baselines. For the former, we asked 11 graduate student volunteers with sufficient training to answer the questions. For the latter, following Wang et al. (2024c), we use GPT-4o to generate random guesses without access to the charts, with the exact prompt detailed in Appendix F.
@@ 4.2 Experimental Results
Table 2 demonstrates the performance of 16 MLLMs on MultiChartQA, from which we derive the following observations. There is a notable disparity between closed-source and open-source MLLMs. The best closed-source model, Claude-3.5-Sonnet, achieves an overall accuracy of 75.30%, while the best open-source model, MiniCPM-V2.6, attains 45.62%. The 29.68% absolute accuracy gap underscores the effectiveness of our benchmark in distinguishing model capabilities through comprehensive questioning.
Models perform poorly on sequential reasoning among four categories. The sequential reasoning questions present significant challenges for MLLMs. For example, Claude-3.5-Sonnet reaches only 52.63% accuracy on this task, with GPT-4o following at 37.89%. The best open-source model, MiniCPM-V2.6, achieves only 24.21%, showing a remarkable performance gap between models and humans.
@@
Chart-domain specialized models exhibit lower accuracy. Both MatCha and ChartInstruct-LLama2, models specifically adapted for chart-related tasks, perform significantly worse than general-purpose models. This may be attributed to the smaller training scale of chart-specific models compared to general-purpose models, as well as the lack of complex reasoning data in their training corpus.
Figure 4: The accuracy of the 14 MLLMs is evaluated under three settings: the original setting, merged charts, and without Chain-of-Thought reasoning. Most models exhibit a decline in performance when processing merged charts or when answering without conducting CoT.
@@ 4.5 The Impact of Chart Reference
We further analyze the impact of including chart references in the input questions. Chart references are phrases that specify which charts the model should focus on to facilitate information localization, such as how many members of the U.S. House are aged 50-59 the 118th Congress in the third chart? We conduct two sets of experiments: Remove chart references. We sample 254 questions from spanning parallel question answering, comparative, and sequential reasoning tasks. The questions are chosen in a way that removing chart references would not impair their answerability. We then remove the chart references from the input questions and evaluate the model’s performance. According to the results in Table 3, the absence of chart references led to a performance drop in most MLLMs. The breakdown of results by question type is detailed in Table 7. These findings suggest that MLLMs still face challenges with information localization in multi-chart scenarios, often relying on chart references to accurately locate the required information. Only input the specified chart. Next, we input only the specific chart containing the required information and evaluate the model using questions from the direct question-answering category. As illustrated in Table 3, providing only the relevant chart leads to accuracy improvements across all models (see Table 8 for detailed results). Notably, for relatively weaker models such as MiniCPM-Llama3-V2.5, Idefics3-8B, and Idefics2-8B, the accuracy gains are particularly substantial. The improvements are especially pronounced for relatively weaker models, such as MiniCPM-Llama3-V2.5, Idefics3-8B, and Idefics2-8B. These findings suggest that MLLMs are easily distracted by irrelevant charts and face challenges in information localization in multi-image settings. This highlights the unique difficulty posed by MultiChartQA compared to previous single-chart evaluations.

### v2

@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have received significant attention due to their impressive capabilities across a broad set of vision-language tasks, such as image captioning Agrawal et al. (2019); Wang et al. (2023), multimodal dialogue systems Saha et al. (2018); Yin et al. (2024), and visual question answering (VQA) Schwenk et al. (2022); Li et al. (2024b); Jia et al. (2024a). Among their varied applications, the comprehension of visual chart data, such as bar and line charts, holds particular significance Han et al. (2023); Zhang et al. (2024a). Such data is prevalent in real-world scenarios, including academic papers and analytical reports, which makes understanding and reasoning over these charts an essential skill for MLLMs. As a result, several efforts have been made in evaluating and improving model performance on chart-related tasks Masry et al. (2022); Xu et al. (2023); Masry et al. (2024a).
@@
To bridge this gap, we introduce MultiChartQA, a newly collected benchmark designed to evaluate the capabilities of MLLMs on multi-chart scenarios. Our benchmark draws from a diverse set of public resources containing multi-chart articles, including Arxiv, OECD, Our World in Data, Pew Research Center, USAFacts, The World Economic Forum, Data Commons, International Energy Agency, GALLUP, and Gapminder. Using the charts collected from these sources, we design four different question types, each paired with multiple charts from the same article as input:
@@
We evaluate the performance of 20 mainstream MLLMs on MultiChartQA. The results reveal a significant limitation in MLLMs’ ability to process multiple charts, with performance trailing behind humans by a large margin. Furthermore, there is a notable gap between the performance of proprietary models and their open-source counterparts. Further analysis conducted on our benchmark highlights the benefits of using chain-of-thought reasoning, as well as the models’ reliance on chart references for information retrieval. An in-depth error analysis reveals that MLLMs still struggle with visual perception and multi-step reasoning, especially when color or spatial constraints are involved. We believe that MultiChartQA will illuminate the evaluation of multi-chart capabilities in MLLMs and foster advancements in this field in the near future.
@@ 3 MultiChartQA
MultiChartQA is an extensive and demanding benchmark that features real-world charts and complex reasoning questions. To cover charts with various topics and a diverse range of styles, we curate chart sets from a variety of real-world sources (§3.1). Based on four core competencies in chart comprehension, we design four categories of question-answer pairs (§3.2). We detail the evaluation metrics in §3.3. Ultimately, the MultiChartQA benchmark comprises 1,370 charts and 2,000 questions.
@@
- Direct Question 527 (26.35%)
- Parallel Questions 617 (30.85%)
- Comparative Reasoning 581 (29.05%)
- Sequential Reasoning 275 (13.75%)
@@
-Multiple-choice 247 (12.35%)
- Open-ended 1,753 (87.65%)
Unique charts 1,370
Multi-chart sets 500
Average charts 2.74
Source 10
@@
Unique questions 1,646
Unique tokens 4,341
Maximum/Average length 162/36.4
@@
Unique answers 1,109
Unique tokens 1,289
Maximum/Average length 55/3.1
@@ 3.1 Chart Collection
For MultiChartQA, we aim to utilize chart sets that maintain semantic coherence and topic consistency. The charts within the same article or paper naturally fulfill this criterion. We therefore identify ten websites that provide such multi-chart content, namely ArXiv [^1] https://arxiv.org/, OECD [^2] https://www.oecd.org/, Our World in Data [^3] https://ourworldindata.org/, Pew Research Center [^4] https://www.pewresearch.org/, USAFacts [^5] https://usafacts.org/, The World Economic Forum [^6] https://www.weforum.org/, Data Commons [^7] https://datacommons.org/, International Energy Agency [^8] https://tinyurl.com/27ywfv6n, GALLUP [^9] https://www.gallup.com/, and Gapminder [^10] https://www.gapminder.org/tools/. We collect chart sets from these sources, covering diverse topics.
Next, we introduce the details of the collection of these sources. ArXiv is an open-access repository that hosts preprints of research papers across various scientific disciplines. We extracted images from 1,500 publications using the arXiv API [^11] https://info.arxiv.org/help/api/index.html. Since research papers often contain a variety of images beyond charts, such as natural images or diagrams, we employ an off-the-shelf MLLM, LLaVA-1.5-13B Liu et al. (2024a), in a zero-shot prompting approach to classify and filter out non-chart images. We consider multiple charts from the same paper as a chart set. To ensure the charts in one set have strong semantic correlations, we manually review the sets and retain those that share one or more elements, e.g., same entities or categories. In total, we collect 129 sets of charts from Arxiv. International Energy Agency contains a variety of charts in different formats and on diverse topics. We manually select charts on this website and group those with the same topic into a chart set, resulting in the collection of 32 multi-chart sets. Similarly, we conduct manual chart collection from Gapminder website, generating a total of 4 chart sets. As to OECD, Our World in Data, Pew Research Center, USAFacts, The World Economic Forum, Data Commons, and GALLUP, we resort to python crawlers for chart downloading, followed by human double-check. We treat the charts from the same report as a semantically related chart set. In total, we collected 58, 58, 56, 50, 48, 44, and 21 multi-chart sets from the aforementioned sources, respectively. Overall, 500 multi-chart sets are collected, each containing 2-3 semantically related charts. The proportion of each source is shown in Appendix H. Some examples of chart sets from different sources can be found in Appendix G.
@@ 3.2 Question Design
The direct questions evaluate whether the model can accurately identify the relevant chart to answer questions accordingly. We present multiple charts and use terms like “In the second chart” to explicitly specify which chart the answer should pertain to. The design of the questions focuses on two types of chart information: structural information and content information. The former involves details about the chart’s structure, such as the values on the axes, the layout of subplots, and the labels in the legend Wang et al. (2024c). These details are essential for comprehending the overall architecture of the chart and interpreting its organizational form. The latter focuses on the chart’s content, such as extracting specific numbers or phrases based on the given clues. This type of task requires the model to understand the semantics of the chart content to perform precise reasoning and information analysis.
@@
Claude-3.5-Sonnet 77.72 86.36 78.42 81.31 80.92 80.65 72.51 79.46 62.55
GPT-4o 72.46 80.77 71.37 76.87 81.42 74.42 75.83 70.41 53.09
Gemini-1.5-Pro 66.83 77.97 71.78 73.84 73.69 62.21 63.03 65.81 48.73
GPT-4o-mini 59.27 72.38 60.58 61.62 63.48 56.91 49.76 64.59 41.45
@@
Qwen2-VL-72B 65.67 81.47 72.61 60.61 68.87 64.98 52.13 73.11 44.36
MiniCPM-V2.6 53.13 69.93 56.85 47.98 54.47 48.85 29.38 66.62 37.82
LLaVA-OV-72B 53.09 67.48 62.24 52.63 62.34 55.30 28.91 58.24 32.36
InternVL2-26B 50.54 65.03 57.26 46.57 55.32 52.30 33.18 55.00 33.82
Qwen2-VL-7B 47.03 64.69 61.83 40.00 48.16 50.00 16.59 60.27 22.18
InternVL-V1.5 46.42 60.14 56.02 38.99 48.79 46.77 27.01 55.68 28.36
LLaVA-OV-7B 36.12 50.70 42.32 29.70 36.10 39.17 12.32 51.76 14.18
MiniCPM-V2.5 33.27 45.45 22.82 30.10 36.10 32.03 20.38 45.81 23.27
DeepSeek-VL-7B 29.42 34.27 32.78 23.23 27.87 28.57 13.27 43.11 21.09
Eagle-X5-13B 28.07 34.97 38.17 20.30 26.95 25.81 12.80 37.97 17.82
LLaVA-V1.6-7B 26.35 26.57 24.90 23.13 27.59 31.34 9.95 37.57 21.82
Idefics3-8B 13.48 11.54 12.86 9.19 12.98 17.74 2.37 24.46 9.45
Idefics2-8B 11.87 7.34 14.11 8.89 10.92 11.06 6.64 22.43 7.64
@@
ChartGemma 11.91 6.99 13.28 5.76 9.22 10.83 6.64 22.84 12.00
TinyChart-3B 10.17 6.99 7.47 7.88 9.79 7.83 6.16 20.68 8.36
MatCha 6.83 7.34 8.71 3.13 2.27 5.07 2.37 16.22 2.91
Table 2: Evaluation results on MultiChartQA. Bold values indicate the best performance within each category: closed-source or open-source models. Question categories and model names are abbreviated due to space limits. Struct.: Structure, Cont.: Content, Seq.: Sequential, LLaVA-OV-72B: LLaVA-OneVision-72B, LLaVA-OV-7B: LLaVA-OneVision-7B
@@ 3.3 Evaluation Metric
We add explicit instructions for each question to control the output format, enabling efficient extraction of answers. The model outputs are parsed and answers are extracted using scripts [^12] https://github.com/MMMU-Benchmark/MMMU.. We then evaluate numerical answers using relaxed accuracy Methani et al. (2020); Masry et al. (2022) and non-numeric answers using exact match to determine whether the extracted answer aligns with the ground truth. For questions that consist of multiple sub-questions, i.e., parallel questions, we calculate accuracy based on the proportion of correctly answered sub-questions, ensuring a fair and fine-grained evaluation.
@@ 4 Experiments
Figure 4: The accuracy of the 18 MLLMs is evaluated under three settings: the original setting, merged charts, and without Chain-of-Thought reasoning. Most models exhibit a decline in performance when processing merged charts or when answering without conducting CoT.
@@ 4.1 Experimental Setup
We evaluate 20 mainstream MLLMs on our benchmark, comprising 4 closed-source models, 13 open-source general-purpose models, and 3 chart-domain specialized models. The closed-source models include Claude-3.5-Sonnet Anthropic (2024), GPT-4o OpenAI (2024), GPT-4o-mini OpenAI (2024), and Gemini-1.5-Pro Reid et al. (2024). Among the open-source general-purpose models, we evaluate Qwen2-VL-72B Wang et al. (2024b), Qwen2-VL-7B Wang et al. (2024b), MiniCPM-V2.6 Yao et al. (2024), MiniCPM-V2.5 Yao et al. (2024), LLaVA-OneVision-72B Li et al. (2024a), LLaVA-OneVision-7B Li et al. (2024a), LLaVA-V1.6-7B Liu et al. (2024b), InternVL2-26B Chen et al. (2024), InternVL-V1.5 Chen et al. (2024), DeepSeek-VL-7B Lu et al. (2024), Eagle-X5-13B Shi et al. (2024), Idefics3-8B Laurençon et al. (2024a), and Idefics2-8B Laurençon et al. (2024b). For the chart-model, we evaluate ChartGemma Masry et al. (2024b), TinyChart-3B Zhang et al. (2024b) and MatCha Liu et al. (2022). For reference, we include human performance and random choice baselines. For the former, we asked 11 graduate student volunteers with sufficient training to answer the questions. For the latter, following Wang et al. (2024c), we use GPT-4o to generate random guesses without access to the charts, with the exact prompt detailed in Appendix F.
@@ 4.2 Experimental Results
Table 2 demonstrates the performance of 20 MLLMs on MultiChartQA, from which we derive the following observations. There is a notable disparity between closed-source and open-source MLLMs. The best closed-source model, Claude-3.5-Sonnet, achieves an overall accuracy of 77.72%. In contrast, most open-source models fall short in performance—except for one, namely Qwen2-VL-72B, which attains an impressive performance of 65.67%. This accuracy gap underscores the effectiveness of our benchmark in distinguishing model capabilities through comprehensive questioning.
Models perform poorly on sequential reasoning among four categories. The sequential reasoning questions present significant challenges for MLLMs. For example, Claude-3.5-Sonnet reaches only 62.55% accuracy on this task, with GPT-4o following at 53.09%. The best open-source model, Qwen2-VL-72B, achieves only 44.36%, showing a remarkable performance gap between models and humans.
@@
Chart-domain specialized models exhibit lower accuracy. All models specifically adapted for chart-related tasks perform significantly worse than general-purpose models. This may be attributed to the smaller training scale of chart-specific models compared to general-purpose models, as well as the lack of complex reasoning data in their training corpus.
@@ 4.5 The Impact of Chart Reference
We further analyze the impact of including chart references in the input questions. Chart references are phrases that specify which charts the model should focus on to facilitate information localization, such as how many members of the U.S. House are aged 50-59 the 118th Congress in the third chart? We conduct two sets of experiments: Remove chart references. We sample 254 questions from spanning parallel question answering, comparative, and sequential reasoning tasks. The questions are chosen in a way that removing chart references would not impair their answerability. We then remove the chart references from the input questions and evaluate the model’s performance. According to the results in Table 3, the absence of chart references led to a performance drop in most MLLMs. The breakdown of results by question type is detailed in Table 7. These findings suggest that MLLMs still face challenges with information localization in multi-chart scenarios, often relying on chart references to accurately locate the required information. Only input the specified chart. Next, we sample 212 questions from the direct question-answering category and evaluate the model using only the specific chart that contains the required information. As illustrated in Table 3, providing only the relevant chart leads to accuracy improvements across all models (see Table 8 for detailed results). Notably, for relatively weaker models such as MiniCPM-Llama3-V2.5, Idefics3-8B, and Idefics2-8B, the accuracy gains are particularly substantial. The improvements are especially pronounced for relatively weaker models, such as MiniCPM-Llama3-V2.5, Idefics3-8B, and Idefics2-8B. These findings suggest that MLLMs are easily distracted by irrelevant charts and face challenges in information localization in multi-image settings. This highlights the unique difficulty posed by MultiChartQA compared to previous single-chart evaluations.